OpenStudy (anonymous):

adult tickets for the show cost $10.50 and each child's ticket cost $7. on sunday the Theater sold 325 tickets for a total revenue of $2982. how many of each ticket did the theater sell on sunday?

OpenStudy (radar):

The first two lines of the above post are good. the third line does not jive.

OpenStudy (anonymous):

ah right i originally used a and c as the cost of each ticket

OpenStudy (anonymous):

then edited the first 2 lines

OpenStudy (radar):

O.K. vg, I did like the way it started out. Just thought the 3rd line would of been a+c=325.

OpenStudy (radar):

It would of went this way. Let c = number of childrens tickets sold Let a = number of adult tickets sold. a+c=325 You now have one equation with two unknowns. You are given cost information lets use that to get another equation 10.50a = amount of money made on adult tickets 7c= amount of money made on children tickets (in dollars of course) $10.50a + $7.00c = $2982 Now you have the second equation with two unknowns. Here are your two equations: a+c=325 $10.50a + $7.00c = $2982 You can use the substitution method to solve it. Did you follow with understanding.

OpenStudy (radar):

Any further questions before I go?